Given

"A=\left(\begin{array}{rr}-2&2\\2&-4\end{array}\right)\quad \textrm{and} \quad B=\left(\begin{array}{rr}2&3\\2&-1\end{array}\right)."


Select the option(s) below which represent "\left(AB^{-1}\right)^{-1}."

  1. "\displaystyle \left(\begin{array}{rr}-3.5&-2.5\\-1.5&-0.5\end{array}\right)."
  2. "\displaystyle \left(\begin{array}{rr}3.5&2.5\\1.5&0.5\end{array}\right)."
  3. A-1B
  4. B-1A-1
  5. BA-1




1
Expert's answer
2021-10-05T11:00:24-0400

Let's find "{B^{ - 1}}" :

"\left( {\left. {\begin{matrix}
2&3\\
2&{ - 1}
\end{matrix}} \right|\begin{matrix}
1&0\\
0&1
\end{matrix}} \right)\begin{matrix}
{}\\
{ - I}
\end{matrix} \Leftrightarrow \left( {\left. {\begin{matrix}
2&3\\
0&{ - 4}
\end{matrix}} \right|\begin{matrix}
1&0\\
{ - 1}&1
\end{matrix}} \right) \Leftrightarrow \left( {\left. {\begin{matrix}
2&3\\
0&1
\end{matrix}} \right|\begin{matrix}
1&0\\
{\frac{1}{4}}&{ - \frac{1}{4}}
\end{matrix}} \right)\begin{matrix}
{ - 3II}\\
{}
\end{matrix} \Leftrightarrow \left( {\left. {\begin{matrix}
2&0\\
0&1
\end{matrix}} \right|\begin{matrix}
{\frac{1}{4}}&{\frac{3}{4}}\\
{\frac{1}{4}}&{ - \frac{1}{4}}
\end{matrix}} \right) \Leftrightarrow \left( {\left. {\begin{matrix}
1&0\\
0&1
\end{matrix}} \right|\begin{matrix}
{\frac{1}{8}}&{\frac{3}{8}}\\
{\frac{1}{4}}&{ - \frac{1}{4}}
\end{matrix}} \right)"

Then

"{B^{ - 1}} = \left( {\begin{matrix}
{\frac{1}{8}}&{\frac{3}{8}}\\
{\frac{1}{4}}&{ - \frac{1}{4}}
\end{matrix}} \right)"

"A{B^{ - 1}} = \left( {\begin{matrix}
{ - 2}&2\\
2&{ - 4}
\end{matrix}} \right)\left( {\begin{matrix}
{\frac{1}{8}}&{\frac{3}{8}}\\
{\frac{1}{4}}&{ - \frac{1}{4}}
\end{matrix}} \right) = \left( {\begin{matrix}
{ - \frac{1}{4} + \frac{1}{2}}&{ - \frac{3}{4} - \frac{1}{2}}\\
{\frac{1}{4} - 1}&{\frac{3}{4} + 1}
\end{matrix}} \right) = \left( {\begin{matrix}
{\frac{1}{4}}&{ - \frac{5}{4}}\\
{ - \frac{3}{4}}&{\frac{7}{4}}
\end{matrix}} \right)"

Let's find "{\left( {A{B^{ - 1}}} \right)^{ - 1}}" :

"{\left( {A{B^{ - 1}}} \right)^{ - 1}} = \left( {\left. {\begin{matrix}
{\frac{1}{4}}&{ - \frac{5}{4}}\\
{ - \frac{3}{4}}&{\frac{7}{4}}
\end{matrix}} \right|\begin{matrix}
1&0\\
0&1
\end{matrix}} \right) \Leftrightarrow \left( {\left. {\begin{matrix}
1&{ - 5}\\
{ - 3}&7
\end{matrix}} \right|\begin{matrix}
4&0\\
0&4
\end{matrix}} \right)\begin{matrix}
{}\\
{ + 3I}
\end{matrix} \Leftrightarrow \left( {\left. {\begin{matrix}
1&{ - 5}\\
0&{ - 8}
\end{matrix}} \right|\begin{matrix}
4&0\\
{12}&4
\end{matrix}} \right) \Leftrightarrow \left( {\left. {\begin{matrix}
1&{ - 5}\\
0&{ - 2}
\end{matrix}} \right|\begin{matrix}
4&0\\
3&1
\end{matrix}} \right)\begin{matrix}
{ - \frac{5}{2}II}\\
{}
\end{matrix} \Leftrightarrow \left( {\left. {\begin{matrix}
1&0\\
0&{ - 2}
\end{matrix}} \right|\begin{matrix}
{ - \frac{7}{2}}&{ - \frac{5}{2}}\\
3&1
\end{matrix}} \right) \Leftrightarrow \left( {\left. {\begin{matrix}
1&0\\
0&1
\end{matrix}} \right|\begin{matrix}
{ - \frac{7}{2}}&{ - \frac{5}{2}}\\
{ - \frac{3}{2}}&{ - \frac{1}{2}}
\end{matrix}} \right)"

Then

"{\left( {A{B^{ - 1}}} \right)^{ - 1}} = \left( {\begin{matrix}
{ - 3.5}&{ - 2.5}\\
{ - 1.5}&{ - 0.5}
\end{matrix}} \right)"

By the properties of the inverse matrix

"{\left( {A{B^{ - 1}}} \right)^{ - 1}} = {\left( {{B^{ - 1}}} \right)^{ - 1}}{A^{ - 1}} = B{A^{ - 1}}"

Answer: 1. "{\left( {A{B^{ - 1}}} \right)^{ - 1}} = \left( {\begin{matrix}
{ - 3.5}&{ - 2.5}\\
{ - 1.5}&{ - 0.5}
\end{matrix}} \right)" , 5. "B{A^{ - 1}}"